# SignCLIP: Connecting Text and Sign Language by Contrastive Learning

### v1

@@ 1 Introduction
SLP spans the tasks of sign language recognition (Adaloglou et al., 2021), translation (De Coster et al., 2023), and production (Rastgoo et al., 2021). Typical datasets, equipped with spoken language text and sign language glosses in addition to videos, that support SLP research are RWTH-PHOENIX-Weather 2014T, in German Sign Language (DGS), introduced by Forster et al. (2014); Camgoz et al. (2018); and CSL-Daily, in Chinese Sign Language (CSL), introduced by Zhou et al. (2021). However, advances on a specific task/dataset/language are often limited and non-transferable to more generic and challenging settings (Müller et al., 2022, 2023) due to the small, domain-specific vocabulary (1,066 and 2,000 signs, respectively) and data size (11 and 23 hours, respectively). Recent sign language corpora of a thousand signing hours have emerged for relatively high-resource sign languages, e.g., BOBSL (Albanie et al., 2021) for British Sign Language (BSL) and YouTube-ASL (Uthus et al., 2024) for American Sign Language (ASL).
@@
Ideally, the visual aspect of sign languages, i.e., larger lexical similarity due to iconicity (Johnston and Schembri, 2007) (illustrated in Figure 1) and smaller grammatical variation (Sandler and Lillo-Martin, 2006) than spoken languages, makes transferring between different sign languages easier than the text of different spoken languages or even scripts. The latter faces unfair tokenization issues (Petrov et al., 2024) and out-of-vocabulary errors. At the same time, unlike discrete text tokens (see the comparison in Table 1), the dense continuous video signal is expensive to process computationally and seems daunting for the above-mentioned self-supervised training approaches.
@@ 2 Background: Sign Language Representation
Representation is a key challenge for SLP. Unlike spoken languages, sign languages have no widely adopted written form. As sign languages are conveyed through the visual-gestural modality, video recording is the most straightforward way to capture them. The final goal of SignCLIP is to represent a sign language video clip by an embedding that aligns with a ubiquitous text encoder like Sentence-BERT (Reimers and Gurevych, 2019). Cheng et al. (2023), a work related to ours, uses a combination of a domain-agnostic and a domain-aware video encoder to address video-to-text retrieval. Generally, end-to-end training on the raw videos is computationally costly, and various intermediate representations alleviate this issue.
@@
A potentially more interpretable and universal way of extracting sign language-related features from videos is human pose estimation (Zheng et al., 2023), for example, using MediaPipe Holistic (Lugaresi et al., 2019; Grishchenko and Bazarevsky, 2020). Each video frame is converted into the location (X, Y, Z) of 543 full body keypoints in a 3D space. However, the immediate applicability of the pose estimation systems for SLP is questionable (Moryossef et al., 2021), and known issues such as the lack of accurate depth information are presented (Holmes et al., 2024). Generally, there is a trade-off between user-friendliness and accuracy for pose estimation tools and SLP researchers often prefer MediaPipe Holistic for the former (Selvaraj et al., 2022) over others like OpenPose (Cao et al., 2019) and AlphaPose (Fang et al., 2022). An even more universal alternative approach to track keypoints named Tracking Everything Everywhere All at Once is proposed by Wang et al. (2023). Sevilla et al. (2024) uses a similar CoTracker (Karaev et al., 2023) model to study sign language prosody. Such models produce smoother signals than traditional pose estimation.
@@ 4 FingerCLIP
As a proof of concept, we first apply this contrastive training approach to Fingerspelling (Battison, 1978; Wilcox, 1992; Brentari and Padden, 2001), a subsystem of sign languages heavily influenced by the surrounding spoken languages. For concepts that do not (yet) have associated signs (names of people, locations, organizations, etc.), sign language users borrow a word of a spoken language by spelling it letter-by-letter with predefined signs for that language’s alphabet. Isolated fingerspelling recognition [^5] Continuous fingerspelling recognition is more complex and is often solved by a CTC loss like speech recognition. can therefore be considered a toy task similar to the MNIST (Deng, 2012) handwritten digits classification task in CV. We name the model FingerCLIP, a mini-version of SignCLIP (§5).
@@
MediaPipe Holistic runs offline on a low-end CPU device. Pose estimation is normalized to a consistent scale by setting the mean width of each person’s shoulders to $1$, and the mid-point to $(0,0)$. The leg values are removed since they are irrelevant to signing. We further augment the data by randomly rotating, shearing, and scaling the poses.
@@
We view fingerspelling understanding as a text-to-video/video-to-text retrieval task. The candidates are ranked for both directions by a dot-product-based similarity score to each text/video query in the latent space. For the test text prompt of each sign, there is possibly more than one correct video (e.g., the same letter signed by different signers) in the test video pool, and they are all considered successful retrieval. We thus evaluate the text-to-video retrieval task by precision@k, i.e., what percent of the top k candidates are correct answers. Each test video query has only one correct text prompt out of the 35 possible prompts. We thus evaluate the video-text retrieval task by recall@k, i.e., the chance to include the only correct answer by taking the top k candidates. precision@1 and recall@1 can be interpreted as the retrieval accuracy. We also add the metric median retrieval rank, i.e., the median value of the rank of the first correct answer in the candidate lists. We present the experimental results in Table 2.
@@ 5 SignCLIP Pretraining
Table 3: Summarization of datasets consisting of relatively short-duration video examples, compared with Spreadthesign and common CV datasets. SignCLIP has been tested with the datasets marked with a checkmark. asl-signs is part of PopSign ASL (Starner et al., 2024), full data not released yet. #signs/#classes for Spreadthesign is marked with an asterisk (*) since the signs of a concept across different sign languages are barely classified as one sign.
@@ 5.1 Spreadthesign Pretraining Dataset
We add the Spreadthesign data scale and that of CLIP and VideoCLIP to Table 3 for comparison. CLIP was trained on 400 million text-image pairs collected from the Internet (500K queries and up to 20K pairs per query); VideoCLIP was pretrained on 1.2 million videos from HowTo100M (each lasts $\sim$6.5 minutes with $\sim$110 clip-text pairs). We additionally add ImageNet (Deng et al., 2009), which has a relatively close scale to Spreadthesign.
@@ 5.2 Training and Evaluation Details
Most implementation details and evaluation protocols in FingerCLIP (§4) are reused for SignCLIP. The text prompts now consist of the text content prepended with a spoken and a sign language tag, inspired by multilingual machine translation in Johnson et al. (2017). For example, the text prompt for signing the phrase “Hello, can I help you?” in ASL is “<en> <ase> Hello, can I help you?”, tagged by the ISO 639-3 language code. Fitting most examples, we limit the context length of the video Transformer $f_{\theta_{v}}$ to 256 tokens, equivalent to a 10-second, 25 FPS video clip; and that of the text Transformer $f_{\theta_{t}}$ to be 64. Both Transformers are initialized with the pretrained bert-base-cased weights and trained on an NVIDIA A100-SXM4-80GB GPU to maximally afford a batch size of 448121212For reference, CLIP was trained with batch size 32,768 and VideoCLIP was trained with batch size 512.. We also measure the training efficiency by the number of parameters and the training time.
For evaluation, we include the same video-to-text metrics used in FingerCLIP and omit text-to-video for simplicity, as the latter correlates with and is more trivial than the former. We additionally test the models with three ASL ISLR datasets in a zero-shot way to evaluate out-of-domain generalization. We perform the following text preprocessing to mitigate the effect of shifted text distribution when building text prompts from the raw gloss labels: (1) lowercasing the glosses; (2) removing the gloss index for different variants of a sign; (3) filtering their test sets by known text labels in Spreadthesign, which reduces the total number of signs/classes.
Starting from a baseline setup that resembles E3 in FingerCLIP, we increase the video Transformer layers from 6 to 12 and add linear multimodal projection layers after temporal pooling. For pose data, we always simplify the face by using the contour keypoints only, resulting in 203 keypoints. We further experiment with the following modifications:
@@
After data normalization, we also employ data augmentation (Boháček and Hrúz, 2022) at training time to improve the models’ robustness, including (1) randomly flipping the poses horizontally; (2) 2D spatial augmentation as done in FingerCLIP; (3) temporal augmentation of the signing speed by linear interpolation between frames; (4) Gaussian noise on keypoints.
@@ 5.3 Experimental Results and Discussion
On the other hand, zero-shot performance on out-of-domain data is deficient. We posit that to reach noticeable performance on out-of-domain data, few-shot learning or fine-tuning (§6.1) is essential given the current scale of pretraining. Nevertheless, starting from E7, we experiment with a few data augmentation techniques to increase data variation since the previous standardization benefits in-domain results but hurts out-of-domain results. We then attempt test-time-only normalization to shift the test distribution of the poses closer to training. As a result, we manage to gradually fight against overfitting and improve the overall zero-shot performance by temporal augmentation (E7.2) and test-time pose flipping if the right hand is not present (E8). The former provides robustness to signing speed change, and the latter is helpful for unseen test examples signed by left-handed signers.
@@ 6.1 Isolated Sign Language Recognition
We provide a comprehensive evaluation for the three ASL ISLR datasets in Table 5. For zero-shot prediction, we follow the optimal setups in E8/E8.1 but skip any text preprocessing on raw glosses for a fairer comparison. For few-shot learning, we randomly sample 10 example videos for each class from training data and use a nonparametric k-nearest neighbors (KNN) algorithm to infer test labels based on pose embedding similarity. We additionally train a supervised logistic regression classifier by scikit-learn on top of the embedded poses, also known as linear probe (Alain and Bengio, 2016). We also train SignCLIP models from scratch or fine-tune them from the E7.2 checkpoint with the three datasets for 25 epochs with batch size 256. Outliers longer than 256 frames are removed.
@@
Table 5: Comprehensive evaluations of SignCLIP on the test set of three ASL ISLR datasets. AS = asl-signs, AC = ASL Citizen, SL = Sem-Lex. The best score SignCLIP achieves on each dataset is in bold. SOTA numbers marked with an asterisk (*) are not directly comparable to ours. SOTA PopSign ASL and SOTA Kaggle are trained with asl-signs but tested on a private test set; SOTA Sem-Lex is tested with a reduced test set of 2,731 classes that are aligned with ASL Citizen/ASL-LEX.
In general, proper zero-shot prediction is hindered by distribution shift on both modalities: (1) for asl-signs dataset, zero-shot prediction is flawed because we only have pose data normalized in an unknown way from Kaggle instead of the raw videos [^13] The authors of PopSign ASL promise to release the full data, then we can do a fairer and more meaningful evaluation.; (2) ASL Citizen uses all upper case glosses while Sem-Lex uses snake case glosses, both unseen during training. For asl-signs, we find pose anonymization eases the domain shift issue in zero-shot (E8.1) and other settings. Besides, pose-based few-shot KNN greatly improves the deficient zero-shot results, bypassing the influence of out-of-domain glosses. Finally, fine-tuning SignCLIP with a pretrained checkpoint, compared to training from scratch, yields closer results to the state-of-the-art (SOTA) reported in the three papers.
@@ 6.2 Sign Language Identification
Sign language identification (Gebre et al., 2013) can be achieved by simply ranking text prompts of different sign languages without actual content, e.g., “<en> <ase>” for ASL. We evaluate the best checkpoint E6.1 for in-domain test data and obtain 0.99 recall@1, which solves the task perfectly without any direct supervision. However, the identification task is ill-defined in that the model can learn to identify signers for particular sign languages.
@@ 6.3 Latent Space Exploration
Iconicity is one of the key motivations for training a multilingual SignCLIP (Figure 1), and now we ask this linguistic question back to the model. We rank a sampled subset of 302 signs from Spreadthesign based on the variance of the pose embeddings across 20+ different sign languages. As a result, the sign for “scorpion” with a universal hook hand shape ranks at the top, and the motivational example “house” sign also ranks high (51/302). On the opposite, the signs for numbers tend to rank low due to the diverse signing styles across languages. We append the full rank in Appendix D.

### v2

@@ 1 Introduction
SLP spans the tasks of sign language recognition (Adaloglou et al., 2021), translation (De Coster et al., 2023), and production (Rastgoo et al., 2021). Typical datasets, equipped with spoken language text and sign language glosses in addition to videos, that support SLP research are RWTH-PHOENIX-Weather 2014T, in German Sign Language (DGS), introduced by Forster et al. (2014); Camgoz et al. (2018); and CSL-Daily, in Chinese Sign Language (CSL), introduced by Zhou et al. (2021). However, advances on a specific task/dataset/language are often limited and non-transferable to more generic and challenging settings (Müller et al., 2022, 2023) due to the small, domain-specific vocabulary (1,066 and 2,000 signs, respectively) and data size (11 and 23 hours, respectively). Recent sign language corpora with a scale of more than a thousand signing hours have emerged for relatively high-resource sign languages, e.g., BOBSL (Albanie et al., 2021) for British Sign Language (BSL) and YouTube-ASL (Uthus et al., 2024) for American Sign Language (ASL), a subset of YouTube-SL-25 (Tanzer and Zhang, 2024). JWSign (Gueuwou et al., 2023) is also a notable corpus that consists of 2,530 hours of Bible translations in 98 sign languages.
@@
Ideally, the visual aspect of sign languages, i.e., lexical similarity due to iconicity (Johnston and Schembri, 2007) (illustrated in Figure 1) makes transferring between different sign languages easier than the text of different spoken languages or even scripts. The latter faces unfair tokenization issues (Petrov et al., 2024) and out-of-vocabulary errors. At the same time, unlike discrete text tokens (see the comparison in Table 1), the dense continuous video signal is expensive to process computationally and seems daunting for the above-mentioned self-supervised training approaches.
@@ 2 Background: Sign Language Representation
Representation is a key challenge for SLP. Unlike spoken languages, sign languages have no widely adopted written form. As sign languages are conveyed through the visual-gestural modality, video recording is the most straightforward way to capture them. The final goal of SignCLIP is to represent a sign language video clip by an embedding that aligns with a ubiquitous text encoder like Sentence-BERT (Reimers and Gurevych, 2019). Generally, end-to-end training on the raw videos is computationally costly, and various intermediate representations alleviate this issue.
@@
A potentially more interpretable and universal way of extracting sign language-related features from videos is human pose estimation (Zheng et al., 2023a), for example, using MediaPipe Holistic (Lugaresi et al., 2019; Grishchenko and Bazarevsky, 2020). Each video frame is converted into the location (X, Y, Z) of 543 full-body keypoints in a 3D space. However, the immediate applicability of the pose estimation systems for SLP is questionable (Moryossef et al., 2021), and known issues such as the lack of accurate depth information are presented (Holmes et al., 2024). Generally, there is a trade-off between user-friendliness and accuracy for pose estimation tools and SLP researchers often prefer MediaPipe Holistic for the former (Selvaraj et al., 2022) over others like OpenPose (Cao et al., 2019) and AlphaPose (Fang et al., 2022). An even more universal alternative approach to track keypoints named Tracking Everything Everywhere All at Once is proposed by Wang et al. (2023). Sevilla et al. (2024) uses a similar CoTracker (Karaev et al., 2023) model to study sign language prosody. Such models produce smoother signals than traditional pose estimation.
@@ 4 FingerCLIP
As a proof of concept, we first apply this contrastive training approach to Fingerspelling (Battison, 1978; Wilcox, 1992; Brentari and Padden, 2001), a subsystem of sign languages heavily influenced by the surrounding spoken languages. For concepts that do not (yet) have associated signs (names of people, locations, organizations, etc.), sign language users borrow a word of a spoken language by spelling it letter-by-letter with predefined signs for that language’s alphabet. Isolated fingerspelling recognition [^5] Continuous fingerspelling recognition is more complex and is often solved by a CTC loss like speech recognition. For more thorough research on fingerspelling, we recommend the ChicagoFSWild+ dataset and Google’s ASL fingerspelling recognition competition on Kaggle. can therefore be considered a toy task similar to the MNIST (Deng, 2012) handwritten digits classification task in CV. We name the model FingerCLIP, a mini-version of SignCLIP (§5).
@@
MediaPipe Holistic runs offline on a low-end CPU device. Pose estimation is normalized to a consistent scale by setting the mean width of each person’s shoulders to $1$, and the mid-point to $(0,0)$. The leg values are removed since they are irrelevant to signing. We further augment the data by randomly rotating, shearing, and scaling the poses [^8] The Pose library implements related operations..
@@
We view fingerspelling understanding as a text-to-video/video-to-text retrieval task. The candidates are ranked for both directions by a dot-product-based similarity score to each text/video query in the latent space. For the test text prompt of each sign, there is possibly more than one correct video (e.g., the same letter signed by different signers) in the test video pool, and they are all considered successful retrieval. We thus evaluate the text-to-video retrieval task by precision@k, i.e., what percent of the top k candidates are correct answers. Each test video query has only one correct text prompt out of the 35 possible prompts. We thus evaluate the video-text retrieval task by recall@k, i.e., the chance to include the only correct answer by taking the top k candidates. precision@1 and recall@1 can be interpreted as the retrieval accuracy, and in our video-to-text scenario where there is only one relevant text item, recall@k also equals top-n accuracy. We keep using recall@k instead of accuracy in this paper for genericity. We also add the metric median retrieval rank, i.e., the median value of the rank of the first correct answer in the candidate lists. We present the experimental results in Table 2.
@@ 5 SignCLIP Pretraining
Table 3: Summarization of datasets consisting of relatively short-duration video examples, compared with Spreadthesign and common CV datasets. SignCLIP has been tested with the datasets marked with a checkmark. asl-signs is a subset of PopSign ASL v1.0. #signs/#classes for Spreadthesign is marked with an asterisk (*) since the signs of a concept across different sign languages are barely classified as one sign.
@@ 5.1 Spreadthesign Pretraining Dataset
We add the Spreadthesign data scale and that of CLIP and VideoCLIP to Table 3 for comparison. CLIP was trained on 400 million text-image pairs collected from the Internet (500K queries and up to 20K pairs per query); VideoCLIP was pretrained on 1.2 million videos from HowTo100M (each lasts $\sim$6.5 minutes with $\sim$110 clip-text pairs). We additionally add ImageNet (Deng et al., 2009), which has a relatively closer scale to Spreadthesign.
@@ 5.2 Training and Evaluation Details
Most implementation details and evaluation protocols in FingerCLIP (§4) are reused for SignCLIP. The text prompts now consist of the text content prepended with a spoken and a sign language tag, inspired by multilingual machine translation in Johnson et al. (2017). For example, the text prompt for signing the phrase “Hello, can I help you?” in ASL is “<en> <ase> Hello, can I help you?”, tagged by the ISO 639-3 language code. Fitting most examples, we limit the context length of the video Transformer $f_{\theta_{v}}$ to 256 tokens, equivalent to a 10-second, 25 FPS video clip; and that of the text Transformer $f_{\theta_{t}}$ to be 64. Both Transformers are initialized with the pretrained bert-base-cased weights and trained on an NVIDIA A100-SXM4-80GB GPU to maximally afford a batch size of 448131313For reference, CLIP was trained with batch size 32,768 and VideoCLIP was trained with batch size 512.. We also measure the training efficiency by the number of parameters and the training time.
For evaluation, we include the same video-to-text metrics used in FingerCLIP and omit text-to-video for simplicity, as the latter correlates with and is more trivial than the former. We additionally test the models with three recent ASL ISLR datasets in a zero-shot way to evaluate out-of-domain generalization. We perform the following text preprocessing to mitigate the effect of shifted text distribution when building text prompts from the raw gloss labels: (1) lowercasing the glosses; (2) removing the gloss index for different variants of a sign [^14] This operation might not be desired if the objective is sign classification as different variants of a sign present different visual forms. The main goal here is to test how well the sign embeddings align to text semantically regardless of the form.; (3) filtering their test sets by known text labels in Spreadthesign, which reduces the total number of signs/classes from the original test datasets.
Starting from a baseline setup that resembles E3 in FingerCLIP, we increase the video Transformer layers from 6 to 12 and try linear multimodal projection layers after temporal pooling. For pose data, we always simplify the face by using the contour keypoints only, resulting in 203 keypoints. We further experiment with the following modifications:
@@
After data normalization, we also employ pose-based data augmentation inspired by Boháček and Hrúz (2022) at training time to improve the models’ robustness, including (1) randomly flipping the poses horizontally; (2) 2D spatial augmentation as done in FingerCLIP; (3) temporal augmentation of the signing speed by linear interpolation between frames; (4) Gaussian noise on keypoints.
@@ 5.3 Experimental Results and Discussion
On the other hand, zero-shot performance on out-of-domain data is deficient. We posit that to reach noticeable performance on out-of-domain data, few-shot learning or fine-tuning (§6.1) is essential given the current scale of pretraining. Nevertheless, starting from E7, we experiment with a few data augmentation techniques to increase data variation given that in-domain standardization benefits in-domain results but hurts out-of-domain results. We then attempt test-time-only normalization to shift the test distribution of the poses closer to training. As a result, we manage to gradually fight against overfitting to the pretraining dataset and improve the overall zero-shot performance by temporal augmentation (E7.2) and test-time pose flipping if the right hand is not present (E8). The former provides robustness to signing speed change in videos, and the latter is helpful for unseen test examples signed by left-handed signers.
@@ 6.1 Isolated Sign Language Recognition
We provide a comprehensive evaluation for ASL ISLR datasets in Table 5. For zero-shot prediction, we follow the optimal setups in E8/E8.1 but skip any text preprocessing on the raw glosses for a fairer comparison. For few-shot learning, we randomly sample 10 example videos for each class from training data and use a nonparametric k-nearest neighbors (KNN) algorithm to infer test labels based on pose embedding similarity ($n\_neighbors=\#classes$). We additionally train a supervised logistic regression classifier with all default settings offered by scikit-learn on top of the embedded poses, also known as linear probe (Alain and Bengio, 2016). We also train SignCLIP models from scratch or fine-tune them from the E7.2 checkpoint with the specific downstream datasets for 25 epochs with batch size 256. Outliers longer than 256 frames are removed.
In general, proper zero-shot prediction is hindered by distribution shift on both modalities: (1) for asl-signs dataset, zero-shot prediction is flawed because we only have pose data normalized in an unknown way from Kaggle instead of the raw videos; (2) PopSign ASL, the superset of asl-signs with raw videos, is released later and the evaluation on it (category game only) is therefore superior and more meaningful; (3) ASL Citizen uses all upper case glosses and Sem-Lex uses snake case glosses, both unseen for the text encoder during training.
@@
PS SignCLIP (E8) zero-shot 0.03 0.10 0.16 62/250
SignCLIP (E8) 10-shot 0.04 0.14 0.22 44
SignCLIP (E8) + linear 0.31 0.58 0.69 4
SignCLIP train from scratch 0.83 0.97 0.99 1
SignCLIP fine-tuned 0.84 0.97 0.98 1
SignCLIP fine-tuned + linear 0.85 0.97 0.98 1
SOTA PopSign ASL 0.84
@@
Table 5: Comprehensive evaluations of SignCLIP on the test set of ASL ISLR datasets. For ISLR, recall@k equals top-n accuracy. AS = asl-signs, PS = PopSign (the superset of AS), AC = ASL Citizen, SL = Sem-Lex. The best score SignCLIP achieves on each dataset is in bold. SOTA numbers marked with an asterisk (*) are not directly comparable to ours. SOTA Kaggle is trained with asl-signs but tested on a private test set; SOTA Sem-Lex is tested with a reduced test set of 2,731 classes that are aligned with ASL Citizen/ASL-LEX and is thus considered an easier objective.
For asl-signs, we find pose anonymization eases the domain shift issue in zero-shot (E8.1) and other settings. Besides, pose-based few-shot KNN greatly improves the deficient zero-shot results, bypassing the influence of out-of-domain glosses. Finally, fine-tuning SignCLIP with a pretrained checkpoint, compared to training from scratch on the target dataset, yields more competitive or even better results than the state-of-the-art (SOTA) reported in the previous literature.
@@ 6.2 Sign Language Identification
Sign language identification (Gebre et al., 2013) can be achieved by simply ranking text prompts of different sign languages without actual content, e.g., “<en> <ase>” for ASL. We evaluate the best checkpoint E6.1 for in-domain test data and obtain 0.99 recall@1, which solves the task perfectly without any direct supervision. However, the identification task is ill-defined in that the model can learn to identify signers for particular sign languages. This task is considered a simplified version of in-domain video-to-text retrieval, as it eliminates the need to simultaneously distinguish language and concept.
@@ 6.3 Latent Space Exploration
Iconicity is one of the key motivations for training a multilingual SignCLIP (Figure 1), and now we ask this linguistic question back to the model. We rank a sampled subset of 302 signs from Spreadthesign based on the variance of the pose embeddings across 20+ different sign languages. As a result, the sign for “scorpion” with a universal hook hand shape ranks at the top, and the motivational example, i.e., the “house” sign, also ranks high (51/302). Conversely, the signs for numbers rank low due to the diverse signing styles across languages. We append the full rank in Appendix D.
7 Related Work
Contrastive learning has been increasingly used in contemporary SLP work. We discuss some related work in this session, which aims mainly at improving performance on dedicated tasks and datasets. Therefore, their goals are slightly different than SignCLIP, a universal pretrained sign language embedding model aligned to spoken language text.
Gan et al. (2023); Zheng et al. (2023b); Ye et al. (2024) focus on the small-scale RWTH-PHOENIX-Weather 2014T and CSL-Daily datasets. Gan et al. (2023) designs a visual contrastive loss and a semantic contrastive loss to tackle the CTC spike phenomenon in the Continuous Sign Language Recognition (CSLR) task and the exposure bias issue in the Sign Language Translation (SLT) task. Zheng et al. (2023b) proposes an explicit contrastive cross-modal alignment between video frames and glosses in combination with an implicit cross-modal alignment by a variational autoencoder (VAE) for CSLR. Ye et al. (2024) identifies a visual representation density issue in SLT and introduces a frame-wise contrastive learning strategy to alleviate the issue and improve SLT.
Wong et al. (2023) has a similar sign embedding idea as SignCLIP, namely Learnt Contrastive Concept, that aligns with word embeddings of the linguistic labels for sign video and incorporates this into ISLR pipelines for the WLASL and BOBSL datasets. Raude et al. (2024), on the other hand, tackles CSLR on BOBSL by a multi-task setting, including two contrastive losses for sign-level and sentence-level retrieval, respectively. They show that joint training for CSLR and sign language retrieval is mutually beneficial.
Like our pretraining task framing introduced in §4, Cheng et al. (2023) explicitly formulates text-to-video/video-to-text retrieval as a cross-lingual contrastive learning task. They address the data scarcity issue by combining a domain-agnostic and a domain-aware video encoder and show its effectiveness on How2Sign and PHOENIX-2014T.